How do you solve this?
2log3 X-log3(X+4) = 2

Answer by rothauserc(4718)

2log3 X-log3(X+4) = 2
:
log(3) (x^2 / (x+4)) = 2
:
Note that 2log(3) x = log(3) x^2, also used log definition of division
:
x^2 / (x+4) = 3^2 = 9
:
Note used definition of logarithm
:
x^2 = 9(x+4)
:
x^2 -9x -36 = 0
:
(x+3) * (x-12) = 0
:
******************************************************
x = 12
:
Note the logarithm of a negative number is not defined
******************************************************
: